# ORCID: 0000-0002-0777-2295

### v1

@@ Abstract
Question-and-answer platforms such as Stack Overflow have become an important way for software developers to share and retrieve knowledge. However, reusing poorly understood code can lead to serious problems, such as bugs or security vulnerabilities. To better understand how code comments affect the perceived helpfulness of Stack Overflow answers, we conducted an online experiment simulating a Stack Overflow environment (n=91). The results indicate that both block and inline comments are perceived as significantly more helpful than uncommented source code. Moreover, novices rated code snippets with block comments as more helpful than those with inline comments. Interestingly, other surface features, such as the position of an answer and its answer score, were considered less important. The content of Stack Overflow has been a major source for training large language models. AI-based coding assistants such as GitHub Copilot, which are based on these models, might change the way Stack Overflow is used. However, our findings have implications beyond this specific platform. First, they may help to improve the relevance of community-driven platforms such as Stack Overflow, which provide human advice and explanations of code solutions, complementing AI-based support for software developers. Second, since chat-based AI tools can be prompted to generate code in different ways, knowing which properties influence perceived helpfulness might lead to targeted prompting strategies to generate more readable code snippets.
@@
Code Comprehension Code Comments Stack Overflow Question-and-answer Platforms User Ratings
@@ 1 Introduction
Programming is a constantly evolving process of problem solving that encompasses tasks such as feature implementation, debugging, and system maintenance. In this dynamic field, software developers often turn to a variety of information sources to achieve their goals. The advent of artificial intelligence (AI) has caused a paradigm shift in programming practices. AI-powered tools have the ability to generate code snippets from user input, streamlining the prototyping process for developers, and scan code for inefficiencies and bugs, while predictive coding and auto-completion features suggest code segments. However, the continued relevance of community-driven platforms such as Stack Overflow, even in the face of technological advances, is underscored by a modest 2.6% (Xue et al., 2023) decline in question activity on the site since the introduction of ChatGPT. It highlights the importance of platforms that provide human advice on programming issues. Stack Overflow, with its large pool of experienced developers, offers an environment in which detailed, community-validated, and context-specific advice thrives. The platform’s peer-review system, manifested through user upvotes and downvotes, is built to ensure that the most accurate and helpful answers are elevated, maintaining a standard of quality and reliability. The diversity of answers to a single question, including their comments—each providing a unique perspective or solution—often yields more comprehensive insights than the solutions typically generated by AI. However, it is also worth noting that there is an active discussion in the development community on if and why Stack Overflow is “fading away” [^1] https://news.ycombinator.com/item?id=41364798 and whether LLMs make Stack Overflow irrelevant. [^2] See also: https://blog.pragmaticengineer.com/are-llms-making-stackoverflow-irrelevant/; https://meta.stackoverflow.com/q/422392/1974143
@@
Despite these discussions, code examples are a critical resource that eases the challenges of independent problem solving. Stack Overflow, with its extensive repository of more than 24 million questions, 36 million answers, and 91 million comments (Stack Exchange Inc, 2025), remains the main question-and-answer (Q&A) forum for software developers around the world. As a platform for learning and sharing, Stack Overflow has revolutionized the way developers seek help and provide solutions to programming questions. It acts as a knowledge-sharing hub, making it easier for developers to share their expertise and find answers to their problems. By sharing their personal experiences and anecdotes, contributors to the platform offer practical insights into the application of programming knowledge that may be beyond the scope of AI-generated advice. The site serves as a comprehensive resource for developers seeking help beyond their local communities. Stack Overflow users include those who ask questions, those who answer them, and those who browse existing Q&A posts for help.
@@
Stack Overflow has attracted significant research interest from both the information systems community (Lee et al., 2019; Sha et al., 2022; Bornfeld and Rafaeli, 2019; Aaltonen and Wattal, 2020) and the empirical software engineering community. Many of the previous studies focused on different properties of the content on Stack Overflow such as text, code, and metadata (Treude et al., 2011; Baltes et al., 2018; Wu et al., 2019) and its relationship to other platforms such as GitHub (Vasilescu et al., 2013; Baltes and Diehl, 2019; Zhang et al., 2019; Abdalkareem et al., 2017). However, there is a lack of research that examines factors from the user’s perspective when selecting a code snippet. Answer selection is critical because even if developers find a relevant post that addresses their problem, they still need to choose an answer that includes a code snippet that can solve their specific problem. Prior research, including that of Treude and Robillard (2017), has identified poor comprehensibility of Stack Overflow code snippets as a common problem that complicates their reuse and can lead to serious threats. These risks include copying poorly understood code, which can lead to bugs (Abdalkareem et al., 2017), security vulnerabilities (Van Der Linden et al., 2020), and licensing issues (Baltes and Diehl, 2019).
@@
To address this research gap, this study investigates the factors that Stack Overflow users consider when selecting code snippets and solutions to their programming problems. Specifically, it examines the role of code style features, such as the type of source code comments (block, inline, or none) and surface-level cues, including answer position and user rating score. Improving the understanding of code snippets with comments is particularly relevant as users increasingly rely on Stack Overflow for complex questions, while simpler questions are often handled by AI tools (Xue et al., 2023). The following research questions guided the study:
@@ 2.1 Stack Overflow as a social Q&A platform for software developers
Social media has changed the way software developers communicate online and coordinate software development activities (Storey et al., 2017). Social communities for software developers, such as GitHub, Twitter/X, or Stack Overflow, provide participants with quick access to knowledge and expertise shared by other developers. These platforms enable various activities, including searching for projects hosted by other developers, contributing to projects, hosting one’s own projects, participating in discussions about programming technologies, and following repositories of interest. Together, individuals form a community around their shared interests on social developer platforms (Vasilescu et al., 2013). Technical Q&A platforms such as Stack Overflow have become increasingly important to software developers as a means of knowledge sharing. As one of the most popular Q&A sites in the world, Stack Overflow serves as a community-based platform for collaborative information sharing across different programming languages. Over the years, Stack Overflow has accumulated a vast amount of programming knowledge consisting of code snippets accompanied by natural language explanations. The popularity and significance of Q&A forums as a source of support for software developers to solve their development problems are substantial. The primary function of Stack Overflow is to provide a platform where users can ask questions and receive answers from other users. To improve the quality of both questions and answers, users have the ability to comment on and edit each other’s posts. In addition, users can earn reputation points and badges for their contributions, e.g., if other users upvote their content.
@@
A fundamental issue that stands out in the literature is that code snippets on Stack Overflow can be difficult to understand, making it challending for developers to reuse parts of the code, potentially leading to security issues. Treude and Robillard (2017) conducted a study to investigate the extent to which developers perceive Stack Overflow code snippets as self-explanatory, identifying several issues such as the organization of the code, naming, incompleteness, quality, domain, clutter, and rationale. Van Der Linden et al. (2020) researched how developers use Stack Overflow as a resource and their interactions with code snippets, focusing on security aspects. They found that developers often select code snippets based on additional information, such as response score and acceptance status, without considering the security of the snippet. Reusing code snippets from Stack Overflow is common also among experienced developers. However, on average, software that utilizes such snippets has a higher percentage of bugs after code reuse (Abdalkareem et al., 2017). Baltes and Diehl (2019) found that developers frequently reuse content from Stack Overflow without the required attribution, making it difficult or even impossible to go back to the original post in case problems arise. Many Stack Overflow code snippets are incomprehensible, of low quality, or require extensive customization for effective reuse (Wu et al., 2019). Consequently, the following section examines the factors that contribute to the readability and comprehensibility of source code.
@@ 2.2 Code reading and comprehension
Source code is a set of computer instructions written in text form using a language that can be understood by humans. Developers are often faced with time-sensitive situations that require rapid product development or bug fixes. In such cases, they rely on existing code to solve their problems (Alarcon et al., 2016). Code transparency refers to its readability and reproducibility. In a study by Alarcon et al. (2016) on transparency, programmers emphasized the importance of elements such as organization, style, and architecture. Among these elements, code style, including comments, was particularly emphasized. Code comments are added to improve readability and support future maintenance activities. The usefulness of comments depends on the programming experience of the reader (Alarcon et al., 2016). Program readability, which refers to the process by which developers understand source code, is a critical aspect of software development. The readability of code differs from that of human-readable text or natural languages due to its structured nature and the inclusion of various elements such as design, documentation, and logic (Buse and Weimer, 2009). The higher the readability of the code, the faster and more accurately a programmer can extract critical information from the program text.
@@
The literature on code comprehension has identified different approaches to reading code. Researchers have proposed several cognitive models Fagerholm et al. (2022) of the reading process to explain how software developers approach the process of code comprehension. These models include top-down, bottom-up, as-needed, and integrated comprehension approaches (Dunsmore et al., 2000).
@@
Brooks (1983) has proposed the top-down theory of code comprehension. According to this theory, software developers begin by forming general hypotheses about the purpose of the code. In the next step, developers attempt to verify these hypotheses by scanning the code for specific structures or operations, which Brooks (1983) calls “beacons” that programmers use to recognize known structures and operations in the code. According to Brooks (1983), code and its documentation can be thought of as a collection of beacons, such as comments, variables, or indentation. When these indicators are found, they confirm the presence of expected structures or operations and validate the developer’s initial hypotheses. However, if no indicators are found, the programmer must examine the code more closely and may need to revise or reject his or her hypotheses. Understanding continues in iterative rounds, in which the software developer repeatedly formulates and verifies hypotheses until he/she has a complete understanding of the entire code (Brooks, 1983; Dunsmore et al., 2000).
@@
The bottom-up theory suggests that software developers start by examining small pieces of code, gradually combining them to form higher levels of abstraction, as their initial knowledge is insufficient to identify beacons (Shneiderman, 1977). In this approach, semantic relationships are grouped into chunks that serve as building blocks for higher-level chunks. The process continues as more chunks are constructed and relationships between existing chunks are recognized, leading to the formation of larger, higher-level chunks. This iterative process helps the software developer develop a general hypothesis about the purpose of the program (Shneiderman, 1977; Dunsmore et al., 2000).
@@
The integrated model of code comprehension combines elements of top-down and bottom-up strategies and suggests that software developers alternate between these strategies depending on the specific context or their familiarity with the code (Von Mayrhauser and Vans, 1995). When the code is familiar, developers tend to use a top-down approach. In contrast, in scenarios where the code is unfamiliar, such as when reading code snippets on Stack Overflow, the bottom-up approach is more often used (Dunsmore et al., 2000).
@@
Beyond code style, code comments, which are the main focus of this paper, are particularly relevant for code comprehension. Pascarella and Bacchelli (2017) have highlighted the importance of source code comments as valuable information for software developers. Comments help improve the readability, comprehensibility, and maintainability of the code (Pascarella and Bacchelli, 2017). Source code and comments are the two most important artifacts for understanding a system (de Souza et al., 2005).
@@
Regarding code comments, Jabrayilzade et al. (2021) developed a taxonomy of 11 inline comment smells based on a multivocal literature review and 899 inline comments from three open-source Java projects that the authors manually labeled. The most common smells were misleading comments that “did not accurately represent what the code does” and obvious comments that “restate what the code does in an obvious manner” (Jabrayilzade et al., 2021). Based on an online survey with 102 participants, Huang et al. (2023) found that inline comments are written more frequently than block comments. In terms of perceived helpfulness, the authors were unable to find a considerable difference. In terms of automatic comment generation, they found that existing models perform better for block comments. Finally, Wang et al. (2024) found that TODO comments in open-source projects are often of low quality.
@@ 2.3 Summary
Although there is a considerable body of knowledge on different properties of Stack Overflow posts (see Section 2.1) and code comprehension in general (see Section 2.2), the impact of different ways of presenting and documenting source code in the typically short code snippets on Stack Overflow has not yet been studied. Stack Overflow’s fragmented code-snippet-based knowledge sharing approach differs from traditional code comprehension, which usually spans larger code bases, potentially even large industrial software projects. For example, results from whole open-source projects regarding the differences between inline and block comments do not necessarily generalize to individual code snippets in Stack Overflow posts. In the specific context of Stack Overflow, users’ assessment of the perceived helpfulness of code snippets can influence their decision to reuse. Better understanding this perceived helpfulness has implications beyond code on Stack Overflow, as AI-based assistants also offer the option to generate multiple solution proposals or re-generate solutions that users are not satisfied with.
@@ 3 Hypotheses
Furthermore, we argue that block source code comments have the potential to provide novice software developers with the additional information they need compared to shorter inline comments. In contrast, experienced programmers tend to prefer inline source comments, which are typically shorter and cause less disruption to the code-reading process. Based on these considerations, we formulated the following hypothesis:
@@ 3.4 Answer score (RQ4)
The Stack Overflow platform uses a scoring system for both questions and answers, allowing users to vote on the helpfulness of a question or answer. Several studies have looked at different aspects of the Stack Overflow community and platform, such as the accepted answer score, the scoring system, and the user reputation. For example, studies have previously examined the correlation between high-scoring answers and accepted answer scores (Abdalkareem et al., 2017). Gantayat et al. (2015) studied Stack Overflow posts and explored the interaction between accepted answers and upvotes. Their results showed that accepting an answer led the community to vote for that answer instead of an alternative answer with a similar or higher score prior to acceptance. Van Der Linden et al. (2020) conducted the first large-scale between-groups experiment and observed that answer details, called surface features, have a significant impact on the selection of both secure and insecure code snippets. The observed positive effect of the answer score on the acceptance and selection of code snippets also has its drawbacks. Due to Stack Overflow’s decentralized structure, its voting and reputation system may be susceptible to manipulation (Chen et al., 2023). Answer details, such as the ‘accepted answer’ mark, user reputation scores, and answer scores, have also been found to mislead users into accepting insecure coding advice, inadvertently promoting vulnerable code (Meng et al., 2018). Building on previous studies that highlight the importance of answer scores in code snippet selection, the following hypothesis was formulated regarding the perceived helpfulness of code snippets:

### v2

@@ Abstract
Question-and-answer platforms such as Stack Overflow are an important way for software developers to share and retrieve knowledge. However, reusing poorly understood code can lead to serious problems, such as bugs or security vulnerabilities. To better understand how code comments affect the perceived helpfulness of Stack Overflow answers, we conducted an online experiment simulating a Stack Overflow environment (n=91). The results indicate that both block and inline comments are perceived as significantly more helpful than uncommented source code. Moreover, novices rated code snippets with block comments as more helpful than those with inline comments. Interestingly, other surface features, such as the position of an answer and its answer score, were considered less important. Moreover, the content of Stack Overflow has been a major source for training large language models. AI-based coding assistants such as GitHub Copilot, which are based on these models, are changing the way Stack Overflow is used. However, our findings have implications beyond Stack Overflow. First, they may help to improve the relevance also of other community-driven platforms, which provide human advice and explanations of code solutions, complementing AI-based support for software developers. Second, since chat-based AI tools can be prompted to generate code in different ways, knowing which properties influence perceived helpfulness can lead to more targeted prompting strategies to generate readable code snippets.
@@
Code Comprehension Code Comments Stack Overflow Question-and-Answer Platforms
@@ 1 Introduction
Programming is a constantly evolving process of problem solving that encompasses tasks such as feature implementation, debugging, and system maintenance. In this dynamic field, software developers often turn to a variety of information sources to achieve their goals. The advent of artificial intelligence (AI) has caused a paradigm shift in programming practices. AI-powered tools have the ability to generate code snippets from user input, streamlining the prototyping process for developers, and scan code for inefficiencies and bugs, while predictive coding and auto-completion features suggest code segments. However, the continued relevance of community-driven platforms such as Stack Overflow, even in the face of technological advances, is underscored by a modest 2.6% (Xue et al., 2023) decline in question activity on the site since the introduction of ChatGPT. It highlights the importance of platforms that provide human advice on programming issues. Stack Overflow, with its large pool of experienced developers, offers an environment in which detailed, community-validated, and context-specific advice thrives. The platform’s peer-review system, manifested through user upvotes and downvotes, is built to ensure that the most accurate and helpful answers are elevated, maintaining a standard of quality and reliability. The diversity of answers to a single question, including their comments—each providing a unique perspective or solution—often yields more comprehensive insights than the solutions typically generated by AI. However, it is also worth noting that there is an active discussion in the community on if and why Stack Overflow is “fading away” [^1] https://news.ycombinator.com/item?id=41364798 and whether LLMs make Stack Overflow irrelevant. [^2] See also: https://blog.pragmaticengineer.com/are-llms-making-stackoverflow-irrelevant/; https://meta.stackoverflow.com/q/422392/1974143
@@
Despite these discussions, code examples are a critical resource that eases the challenges of independent problem solving. Stack Overflow, with its extensive repository of more than 24 million questions, 36 million answers, and 91 million comments (Stack Exchange Inc, 2025), remains the main question-and-answer (Q&A) forum for software developers around the world. As a platform for learning and sharing, Stack Overflow has revolutionized the way developers seek help and provide solutions to programming questions. It acts as a knowledge-sharing hub, making it easier for developers to share their expertise and find answers to their problems. By sharing their personal experiences and anecdotes, contributors to the platform offer practical insights into the application of programming knowledge that is more personal than AI-generated advice. The site serves as a comprehensive resource for developers seeking help beyond their local communities. Stack Overflow users include those who ask questions, those who answer them, and those who browse existing Q&A posts for help.
@@
Stack Overflow has attracted significant research interest from both the information systems community (Lee et al., 2019; Sha et al., 2022; Bornfeld and Rafaeli, 2019; Aaltonen and Wattal, 2020) and the empirical software engineering community. Many of the previous studies focused on different properties of the content on Stack Overflow such as text, code, and metadata (Treude et al., 2011; Baltes et al., 2018; Wu et al., 2019) and its relationship to other platforms such as GitHub (Vasilescu et al., 2013; Baltes and Diehl, 2019; Zhang et al., 2019; Abdalkareem et al., 2017). However, there is a lack of research that examines factors from the user’s perspective when selecting a code snippet. Answer selection is critical because even if developers find a relevant post that addresses their problem, they still need to choose an answer that includes a code snippet that can solve their specific problem. Prior research, including that of Treude and Robillard (2017), has identified poor comprehensibility of Stack Overflow code snippets as a common problem that complicates their reuse. Potential risks include copying poorly understood code, which can lead to bugs (Abdalkareem et al., 2017), security vulnerabilities (Van Der Linden et al., 2020), and licensing issues (Baltes and Diehl, 2019).
@@
To address this research gap, this study investigates the factors that Stack Overflow users consider when selecting code snippets and solutions to their programming problems. Specifically, it examines the role of code style features, such as the type of source code comments (block, inline, or none) and surface-level cues, including answer position and score. Improving the understanding of code snippets with comments is particularly relevant as users increasingly rely on Stack Overflow for complex questions, while simpler questions are often handled by AI tools (Xue et al., 2023). The following research questions guided the study:
@@ 2.1 Stack Overflow as a social Q&A platform for software developers
Social media has changed the way software developers communicate online and coordinate software development activities (Storey et al., 2017). Social communities for software developers, such as GitHub, Twitter/X, or Stack Overflow, provide participants with quick access to knowledge and expertise shared by other developers. These platforms enable various activities, including searching for projects hosted by other developers, contributing to projects, hosting one’s own projects, participating in discussions about programming technologies, and following repositories of interest. Together, individuals form a community around their shared interests on social developer platforms (Vasilescu et al., 2013). Technical Q&A platforms such as Stack Overflow have become increasingly important to software developers as a means of knowledge sharing. As one of the most popular Q&A sites in the world, Stack Overflow serves as a community-based platform for collaborative information sharing spanning diverse programming languages. Over the years, Stack Overflow has accumulated a vast amount of programming knowledge consisting of code snippets accompanied by natural language explanations. The popularity and significance of Q&A forums as a source of support for software developers to solve their development problems are substantial. The primary function of Stack Overflow is to provide a platform where users can ask questions and receive answers from other users. To improve the quality of both questions and answers, users have the ability to comment on and edit each other’s posts. In addition, users can earn reputation points and badges for their contributions, e.g., if other users upvote their content.
@@
A fundamental issue that stands out in the literature is that code snippets on Stack Overflow can be difficult to understand, making it challenging for developers to reuse parts of the code, potentially leading to security issues. Treude and Robillard (2017) conducted a study to investigate the extent to which developers perceive Stack Overflow code snippets as self-explanatory, identifying several issues such as the organization of the code, naming, incompleteness, quality, domain, clutter, and rationale. Van Der Linden et al. (2020) researched how developers use Stack Overflow as a resource and their interactions with code snippets, focusing on security aspects. They found that developers often select code snippets based on additional information, such as response score and acceptance status, without considering the security of the snippet. Reusing code snippets from Stack Overflow is common also among experienced developers. However, on average, software that utilizes such snippets has a higher percentage of bugs after code reuse (Abdalkareem et al., 2017). Baltes and Diehl (2019) found that developers frequently reuse content from Stack Overflow without the required attribution, making it difficult or even impossible to go back to the original post in case problems arise. Many Stack Overflow code snippets are incomprehensible, of low quality, or require extensive customization for effective reuse (Wu et al., 2019). Consequently, the following section examines the factors that contribute to the readability and comprehensibility of source code.
@@ 2.2 Code reading and comprehension
Source code is a set of computer instructions written in text form using a language that can be understood by humans. Developers often face time-sensitive situations that require rapid product development or bug fixes. In such cases, they rely on existing code to solve their problems (Alarcon et al., 2016). Code transparency refers to its readability and reproducibility. In a study by Alarcon et al. (2016) on transparency, programmers emphasized the importance of elements such as organization, style, and architecture. Among these elements, code style, including comments, was particularly emphasized. Code comments are added to improve readability and support future maintenance activities. The usefulness of comments depends on the programming experience of the reader (Alarcon et al., 2016). Program readability, which refers to the process by which developers understand source code, is a critical aspect of software development. The readability of code differs from that of human-readable text or natural languages due to its structured nature and the inclusion of various elements such as design, documentation, and logic (Buse and Weimer, 2009). The higher the readability of the code, the faster and more accurately a programmer can extract critical information from the program text.
@@
The literature on code comprehension has identified different approaches to reading code. Researchers have proposed several cognitive models (Fagerholm et al., 2022) of the reading process to explain how software developers approach the process of code comprehension. These models include top-down, bottom-up, as-needed, and integrated comprehension approaches (Dunsmore et al., 2000).
@@
Brooks (1983) has proposed the top-down theory of code comprehension. According to this theory, software developers begin by forming general hypotheses about the purpose of the code. In the next step, developers attempt to verify these hypotheses by scanning the code for specific structures or operations, which Brooks (1983) calls “beacons” that programmers use to recognize known structures and operations in the code. According to Brooks (1983), code and its documentation can be thought of as a collection of beacons, such as comments, variables, or indentation. When these indicators are found, they confirm the presence of expected structures or operations and validate the developer’s initial hypotheses. However, if no indicators are found, the programmer must examine the code more closely and may need to revise or reject their hypotheses. Understanding continues in iterative rounds, in which software developers repeatedly formulate and verify hypotheses until they have a complete understanding of the entire code (Brooks, 1983; Dunsmore et al., 2000).
@@
The bottom-up theory suggests that software developers start by examining small pieces of code, gradually combining them to form higher levels of abstraction, as their initial knowledge is insufficient to identify beacons (Shneiderman, 1977). In this approach, semantic relationships are grouped into chunks that serve as building blocks for higher-level chunks. The process continues as more chunks are constructed and relationships between existing chunks are recognized, leading to the formation of larger, higher-level chunks. This iterative process helps the software developer form a general hypothesis about the purpose of the program (Shneiderman, 1977; Dunsmore et al., 2000).
@@
The integrated model of code comprehension combines elements of top-down and bottom-up strategies and suggests that software developers alternate between these strategies depending on the specific context or their familiarity with the code (Von Mayrhauser and Vans, 1995). When the code is familiar, developers tend to use a top-down approach. In contrast, in scenarios where the code is unfamiliar, such as when reading code snippets on Stack Overflow, the bottom-up approach is more frequently used (Dunsmore et al., 2000).
@@
Beyond code style, code comments, which are the main focus of this paper, are particularly relevant for code comprehension. Pascarella and Bacchelli (2017) have highlighted the importance of source code comments as an information for software developers. Comments help improve the readability, comprehensibility, and maintainability of the code (Pascarella and Bacchelli, 2017). Source code and comments are the two most important artifacts for understanding a system (de Souza et al., 2005).
@@
Regarding code comments, Jabrayilzade et al. (2021) developed a taxonomy of 11 inline comment smells based on a multivocal literature review and 899 inline comments from three open-source Java projects that the authors manually labeled. The most common smells were misleading comments that did “not accurately represent what the code does” and obvious comments that “restate what the code does in an obvious manner” (Jabrayilzade et al., 2021). Based on an online survey with 102 participants, Huang et al. (2023) found that inline comments are written more frequently than block comments. Regarding perceived helpfulness, the authors were unable to find a considerable difference. With respect to automatic comment generation, they found that existing models perform better for block comments. Finally, Wang et al. (2024) found that TODO comments in open-source projects are often of low quality.
@@ 2.3 Summary
Although there is a considerable body of knowledge on different properties of Stack Overflow posts (see Section 2.1) and code comprehension in general (see Section 2.2), the impact of different ways of presenting and documenting source code in the typically short code snippets on Stack Overflow has not yet been studied. Stack Overflow’s fragmented code-snippet-based knowledge sharing approach differs from traditional code comprehension, which usually spans larger code bases, potentially even large industrial software projects. For example, results from whole open-source projects regarding the differences between inline and block comments do not necessarily generalize to individual code snippets in Stack Overflow posts. In the specific context of Stack Overflow, users’ assessment of the perceived helpfulness of code snippets can influence their decision to reuse. A better understanding of perceived helpfulness has implications beyond code on Stack Overflow, as AI-based assistants also offer the option to generate multiple solution proposals or re-generate solutions that users are not satisfied with.
@@ 3 Hypotheses
Furthermore, we argue that block source code comments have the potential to provide novice software developers with the additional information they need compared to shorter inline comments. In contrast, experienced programmers tend to prefer inline source comments, which are typically shorter and cause less disruption to the code-reading process. Based on these considerations, we formulate the following hypothesis:
@@ 3.4 Answer score (RQ4)
The Stack Overflow platform uses a scoring system for both questions and answers, allowing users to vote on the helpfulness of a question or answer. Several studies have looked at different aspects of the Stack Overflow community and platform, such as the accepted answer score, the scoring system, and the user reputation. For example, Gantayat et al. (2015) studied Stack Overflow posts and explored the interaction between accepted answers and upvotes. Their results showed that accepting an answer led the community to vote for that answer instead of an alternative answer with a similar or higher score prior to acceptance. Van Der Linden et al. (2020) conducted the first large-scale between-groups experiment and observed that answer details, called surface features, have a significant impact on the selection of both secure and insecure code snippets. The observed positive effect of the answer score on the acceptance and selection of code snippets also has its drawbacks. Due to Stack Overflow’s decentralized structure, its voting and reputation system may be susceptible to manipulation (Chen et al., 2023). Answer details, such as the ‘accepted answer’ mark, user reputation scores, and answer scores, have also been found to mislead users into accepting insecure coding advice, inadvertently promoting vulnerable code (Meng et al., 2018). Building on previous studies that highlight the importance of answer scores in code snippet selection, the following hypothesis was formulated regarding the perceived helpfulness of code snippets: